OpenStudy (anonymous):

ANSWER AND I’LL FAN U AND GIVE U A MEDAL COMBINED INEQUAILITES …. Solve the open sentence. –3 ≤ n + 6 ≤ 7 A.n ≤ –9 and n ≤ 6 B.n ≥ –9 and n ≤ 1 C.n ≤ 3 and n ≤ 7 D.n ≥ 9 and n ≤ 7

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

thnx @tyree20

OpenStudy (anonymous):

youre welcome

OpenStudy (anonymous):

-3 <= n + 6 <= 7 --- subtract 6 -3 - 6 <= n + 6 - 6 <= 7 - 6 -9 <= n <= 1 your right...it is B

OpenStudy (anonymous):

:))) thank yew also...lool ! @kelliegirl33

OpenStudy (anonymous):

I just thought you might want to see how it was done :)

OpenStudy (anonymous):

yurp..! @kelliegirl33 ur good at reading my mind..lool..!..

OpenStudy (anonymous):

lol...